OpenStudy (anonymous):

Solve D = 1/4(e – 3f) for e. You must show all work.

OpenStudy (anonymous):

Starting here: D = 1/4 (e-3f) Multiply both sides by 4: 4D = e-3f Add 3f to both sides: 4D+3f = e Make sense?

OpenStudy (anonymous):

Oh, okay I understand looks easy now!

OpenStudy (anonymous):

You're welcome come back anytime lol :D